Problem 3-16 Comprehensive Problem [LO3-1, LO3-2, LO3-4]

Gold Nest Company of Guandong, China, is a family-owned enterprise that makes birdcages for the South China market. The company sells its birdcages through an extensive network of street vendors who receive commissions on their sales.

The company uses a job-order costing system in which overhead is applied to jobs on the basis of direct labor cost. Its predetermined overhead rate is based on a cost formula that estimated $75,000 of manufacturing overhead for an estimated activity level of $50,000 direct labor dollars. At the beginning of the year, the inventory balances were as follows:

Raw materials $ 11,000
Work in process $

4,500

Finished goods $ 8,600

During the year, the following transactions were completed:

  1. Raw materials purchased on account, $ 164,000.
  2. Raw materials used in production, $146,000 (materials costing $123,000 were charged directly to jobs; the remaining materials were indirect).
  3. Costs for employee services were incurred as follows:
Direct labor $ 153,000
Indirect labor $ 147,000
Sales commissions $ 22,000
Administrative salaries $

44,000

  1. Rent for the year was $18,800 ($13,500 of this amount related to factory operations, and the remainder related to selling and administrative activities).
  2. Utility costs incurred in the factory, $15,000.
  3. Advertising costs incurred, $11,000.
  4. Depreciation recorded on equipment, $22,000. ($17,000 of this amount related to equipment used in factory operations; the remaining $5,000 related to equipment used in selling and administrative activities.)
  5. Record the manufacturing overhead cost applied to jobs.
  6. Goods that had cost $229,000 to manufacture according to their job cost sheets were completed.
  7. Sales for the year (all paid in cash) totaled $514,000. The total cost to manufacture these goods according to their job cost sheets was $219,000.

Required:

1. Prepare journal entries to record the transactions for the year.

2. Prepare T-accounts for each inventory account, Manufacturing Overhead, and Cost of Goods Sold. Post relevant data from your journal entries to these T-accounts (don’t forget to enter the beginning balances in your inventory accounts).

3A. Is Manufacturing Overhead underapplied or overapplied for the year?

3B. Prepare a journal entry to close any balance in the Manufacturing Overhead account to Cost of Goods Sold.

4. Prepare an income statement for the year. All of the information needed for the income statement is available in the journal entries and T-accounts you have prepared.
Answer #1

Part 1

No.

Transaction

General journal

Debit

Credit

1

a

Raw materials

164000

cash

164000

2

b

Work in process

123000

Manufacturing overhead

23000

Raw materials

146000

3

c

Work in process

153000

Manufacturing overhead

147000

Sales commissions expense

22000

Administrative salaries

44000

cash

366000

4

d

Manufacturing overhead

13500

Rent expense

5300

Cash

18800

5

e

Manufacturing overhead

15000

cash

158000

6

f

Advertising expense

11000

Cash

11000

7

g

Manufacturing overhead

17000

Depreciation expense

5000

Accumulated depreciation

22000

8

h

Work in process

229500

Manufacturing overhead

229500

9

i

Finished goods

229000

Work in process

229000

10

j

Cash

514000

Sales

514000

Cost of goods sold

219000

Finished goods

219000

Predetermined overhead rate = Estimated manufacturing overhead cost/Estimated total amount of the allocation base = 75000/50000= 150%

Manufacturing overhead applied = direct labor cost * Predetermined overhead rate = 153000*150% = 229500

Part 2

Raw materials

Beg bal.

11000

(b)

146000

(a)

164000

End. Bal.

29000

Work in process

Beg bal.

4500

(i)

229000

(b)

123000

(c)

153000

(h)

229500

End. Bal.

281000

Finished goods

Beg bal.

8600

(j)

219000

(i)

229000

End. Bal.

18600

Manufacturing overhead

(b)

23000

(h)

229500

(c)

147000

(d)

13500

(e)

15000

(g)

17000

End. Bal.

14000

Cost of goods sold

(j)

219000

Part 3 a

Underapplied.

Manufacturing overhead underapplied = actual manufacturing overhead - manufacturing overhead applied

Part 3 b

General Journal

Debit

Credit

Cost of goods sold

14000

Manufacturing overhead

14000

Part 4

Gold Nest Company

Income Statement

Sales

514000

Cost of goods sold

233000

Gross profit

281000

Selling and administrative expenses

Sales commissions

22000

Administrative salaries

44000

Rent expense

5300

Advertising expense

11000

Depreciation expense

5000

87300

Net operating income

$193700

Cost of goods sold = 219000+14000= 693000